Calculating Departmental Overhead Rates and Applying Overhead to Production

At the beginning of the year, Glaser Company estimated the following:

Assembly
Department
Testing
Department

Total
Overhead $763,000 $740,110 $1,503,110
Direct labor hours 157,000 130,310 287,310
Machine hours 63,750 36,338 100,088

Glaser uses departmental overhead rates. In the assembly department, overhead is applied on the basis of direct labor hours. In the testing department, overhead is applied on the basis of machine hours. Actual data for the month of March are as follows:

Assembly
Department
Testing
Department

Total
Overhead $61,040 $59,209 $120,249
Direct labor hours 12,560 10,425 22,985
Machine hours 5,100 2,907 8,007

Required:

1. Calculate the predetermined overhead rates for the assembly and testing departments. Round your answers to the nearest cent.

Assembly department overhead rate $ per direct labor hour
Testing department overhead rate $ per machine hour

2. Calculate the overhead applied to production in each department for the month of March. If required, round your answers to nearest whole value.

Assembly department $
Testing department $

3. By how much has each department's overhead been overapplied? Underapplied?

Assembly department Overapplied $
Testing department Overapplied $
Answer #1

1) Overhead rate :

Assembly department overhead rate = 763000/157000 = 4.86 per labor hour

Testing department overhead rate = 740110/36338 = $20.37 per machine hour

2) Overhead applied

Assembly department = 12560*4.86 = $61042

Testing department = 2907*20.37 = $59215

3) Calculate following

Assembly department Over applied 61042-61040 = $2
Testing department Over applied 59215-59209 = $6